# Addressing Noise and Efficiency Issues in Graph-Based Machine Learning Models From the Perspective of Adversarial Attack

### v1

@@ IV.5 Efficacy of Resolving Computational Bottleneck
Figure 1 and Table 2 show the eigen-decomposition time of the original full data set and our selected robust node set. It can be observed that by utilizing the selected robust node set, eigendecomposition has been accelerated by a factor of 9 for USPS and 90 for MNIST, respectively. It is expected that the proposed method will be a key enabler for running computationally expensive graph algorithms in scenarios that require extremely fast response times and on devices with limited computational capabilities.
@@
MNIST 6.35 0.07
Figure 1: Eigen-decomposition time of original data set and selected robust nodes .
V Conclusion
In this paper, we have introduced a perspective that views the noisy edges in graphs used by graph algorithms through the lens of adversarial attacks. Building upon this viewpoint, we proposed a method based on robust nodes to mitigate the impact of noisy edges, grounded in adversarial robustness evaluation. Experimental results demonstrate that our approach significantly outperforms state-of-the-art methods for handling noisy edges by a large margin.

### v2

@@ 4.5 Efficacy of Resolving Computational Bottleneck
Figure 1 and Table 2 show the eigen-decomposition time of the original full data set and our selected robust node set. It can be observed that by utilizing the selected robust node set, eigen-decomposition has been accelerated by a factor of 9 for the USPS data set and 90 for the MNIST data set, respectively. It is expected that the proposed method will be a key enabler for running computationally expensive graph-based algorithms in scenarios that require extremely fast response times and on devices with limited computational capabilities.
@@
MNIST(test) 6.35 0.07
Figure 1: Eigen-decomposition time of original node set and robust node set.
4.6 Parameter Discussion
In our method, the construction of $G_{\text{input}}$ and $G_{\text{output}}$ involves the selection of k in the k-NN algorithm. Our preliminary experiments indicate that when k is large enough to adequately capture the underlying structure of the data, the algorithm is not highly sensitive to the parameter setting. This means that the parameter choice is not overly strict. For example, when constructing $G_{\text{input}}$, we can choose k=50, and when constructing $G_{\text{output}}$, we can select k=10, achieving very good results. It is important to note that the k-NN graph is just one common method to capture the underlying structure of the data; it is neither the only nor the necessary approach. Other more advanced graph learning methods can also be used.
5 Conclusion
In this paper, we view the noisy edges in graphs used from the perspective of adversarial attack. Building upon this viewpoint, we proposed a method based on robust nodes to mitigate the impact of noisy edges, grounded in adversarial robustness evaluation. Experimental results from real-world datasets show that our approach significantly boosts the performance of graph-based algorithms and outperforms the state-of-the-art methods in addressing noisy edges by a large margin.